Use the information provided to write the standard form equation of ellipse. Then sketch



the graph.



Foci: (-10, 16), (-10,-8)



Endpoints of major axis: (-10, 17),(-10,-9)

1
Expert's answer
2021-11-25T13:38:04-0500

Vertices: "(-10, 17),(-10,-9)"

Major axis is vertical, length "=17-(-9)=26=2a=>a=13"


"17-13=4=-9-(-13)"

Center: "(-10, 4)"


Foci: "(-10, 16), (-10,-8)"


"2c=16-(-8)=24=>c=12"

"c^2=a^2-b^2=>b^2=13^2-12^2=25=>b=5"

The equation of the ellipse


"\dfrac{(y-4)^2}{169}+\dfrac{(x+10)^2}{25}=1"